1. Identify the conic section, convert to standard form, and then give a rough sketch (a) x? - 9y2 - 2x - 36y - 44 = 0
Answer #1


(9) x² - gy² - 2x - 364 -44=0 x?-gu+1-1-9[y?+4y) - 4+ = 0 (x-1% -9[42+4474-4] - 45-6 Q_y? – 919 +23° +36-4570 (x-3)* - 9(4+2) * ( 1 - 3 ( 1 ) ง 3